How do i calculate the maximum profit of -10x^2 + 1700x - 60000
Answer by stanbon(48568)

How do i calculate the maximum profit of
P(x) = -10x^2 + 1700x - 60000
---
You have a quadratic with a = -10, b = 1700, c = -60000
-------
Max occurs when x = -b/2a = -1700/(2*-10) = 85
-------
P(85) = -10*(85)^2 + 1700*85 - 600000
P(85) = $12,250
===================
Cheers,
Stan H.